# How Far Are We from Believable AI Agents? A Framework for Evaluating the Believability of Human Behavior Simulation

### v1

@@ Abstract
Human behavior simulation of AI agents necessitates the agents to possess a quality of believability, which is crucial as it facilitates users in establishing trust toward the agents and streamlines the fulfillment of the agents’ goal. While recent advancements in Large Language Model (LLM) based agents have improved human behavior simulation, challenges inherent to LLMs (e.g., long context modeling) can undermine their believability. Consequently, evaluating AI agent believability becomes imperative. Unfortunately, prior research often neglects the negative impacts of LLM deficiencies. To address these gaps, we introduce two metrics for assessing LLM-based agent believability: consistency, and robustness, together with a benchmark, SimulateBench, with which, we evaluate the consistency and robustness of agents implemented with popular LLMs. We find that agents (i) struggle to accurately depict character information when presented with lengthy profile inputs; (ii) exhibit vulnerability to profile perturbations, and (iii) are significantly affected by certain key factors that impact their overall believability.  [^1] Code and SimulateBench are public at GitHub
@@ 1 Introduction
AI agents (Wooldridge and Jennings, 1995; Macal and North, 2005) have shown promise to simulate human behavior and social interaction, which can empower applications ranging across prototyping social theories (Aher et al., 2023b; Horton, 2023a; Kosinski, 2023; Kovac et al., 2023), generating synthetic research data (Hämäläinen et al., 2023; Wang et al., 2023b) and building non-player characters (Laird and VanLent, 2001). These applications necessitate the agents’ simulated human behavior to possess a quality of believability : the generated behavior enables others to suspend their disbelief towards the agents (Ortony et al., 2003), which is crucial because it facilitates others in establishing trust toward the agents and streamlines the fulfillment of the agents’ goals in these applications.
@@
Figure 1: An illustrative example of the “Consistency”, and “Robustness”. Consistency measures whether the LLMs’ generated human behavior accurately depicts the identity information; Robustness measures whether the generated human behavior will be influenced by the perturbation in the profile.
@@
Recently, the LLM-based agents (Park et al., 2023) have accelerated the realization of human behavior simulation by providing the LLMs with some identity information through prompting engineering (Liu et al., 2023b). Despite the initial results achieved now, some challenging issues of LLMs may pose a significant threat to the believability of LLM-based agents: (1) current LLMs do not perfectly make use of information in long input contexts (Liu et al., 2023a). This may impede LLM-based agents from capturing crucial information about characters in the long profile, which is usually integrated into the input to prompt agents to simulate humans. The profile may consist of personal data such as identity information, social roles (Wasserman and Faust, 1994), and relationships according to the simulation target. (2) LLMs lack robustness confronted with perturbed inputs (Perez and Ribeiro, 2022). Hence, minimal alterations to the profile input may result in significant fluctuations in the model’s performance within the same scenario, which will undermine the model’s believability. Unfortunately, the profile information usually undergoes dynamic alterations, such as a change in age, because agents are constantly evolving in the environment, which will cause the profile to update. Nevertheless, current studies (Park et al., 2023; Wang et al., 2023c; Shao et al., 2023; Cheng et al., 2023) fail to evaluate the believability along with the two issues. Firstly, the metrics proposed in these works for evaluation are not indicative of the agent’s believability when processing long and diverse profile input because these works only prompt limited identity information to the agents, which is of a concise nature. Secondly, there is little or no mechanism to thoroughly examine the potential impact of perturbations on the believability of LLMs and how these perturbations specifically affect their credibility.
@@
To assess the negative impact brought by LLMs’ deficiency in processing long profile input and lack of robustness, we propose two metrics to evaluate the believability of LLM-based agents, as shown in Figure 1: (1) consistency: to what extent does the LLM-generated human behavior accurately depict the identity information, social roles, and relationships presented in the long profile input? (2) robustness: to what extent do the agents’ behaviors maintain robustness when faced with perturbations in the profile? To measure the consistency and robustness, we introduce SimulateBench: a framework for character data collection and evaluation of agents’ consistency and robustness ability. SimulateBench consists of three parts: the profile descriptive framework, the consistency dataset, and the robustness dataset. The profile descriptive framework is proposed to document the information about a person comprehensively. Based on the framework, we collect character datasets on characters for the agents to simulate. To measure the consistency, we assess whether the agent can correctly answer single-choice questions about the character that the agent is assigned to simulate in the consistency dataset. To measure the robustness, we perturb the profile of the assigned character and compare how the agent’s consistency ability changed faced with the profile perturbation.
@@
Through the SimulateBench, we evaluate the level of believability of agents composed of API-based or open-source LLMs. Our findings show that: (1) agents perform poorly for consistency: they can not accurately depict the information in the long profile input, even if they are equipped with long context size; (2) agents exhibit a lack of robustness when faced with even nuanced profile perturbation. In further studies, we find that demographic factors, information position in the profile, and reasoning prompting will greatly influence the agents’ believability.
@@
In summary, we propose two novel metrics of consistency and robustness to measure agents’ believability. To facilitate the assessment, we introduce the SimulateBench. We hope our work will inspire further research into the believability of agents’ human behavior simulation.
@@ 2 Related Work
2.1 LLM based Agent for Simulation
Recently, LLMs demonstrate intelligence comparable to humans (Rao et al., 2023a; Li et al., 2023; Jiang et al., 2023a; bench authors, 2023; Brown et al., 2020; Touvron et al., 2023; Chowdhery et al., 2022). Many studies endeavor to harness the LLMs as a core brain in agents to simulate human behavior and social interactions (Maes, 1995; Wooldridge and Jennings, 1995; Xi et al., 2023). Park et al. (2022); Gao et al. (2023a) employs agents to simulate genuine human emotions, attitudes, and interaction behaviors in social networks. Park et al. (2023) applies in-context learning to enable the agent to emulate human behaviors in a sandbox world, such as organizing a party. Argyle et al. (2022) shows that the “algorithmic bias” within GPT-3 can be used to emulate response distributions from human subgroups. Horton (2023b) proposes LLMs can be used in behavioral economics experiments to pilot studies. Hämäläinen et al. (2023); Wang et al. (2023b) lets LLM-based agents simulate human behavior in a recommender system to gather user behavior data. In summary, LLM-based agents are being employed to simulate human behaviors and social interactions across several applications.
2.2 Evaluation for simulation of LLM-based Agents
Simulation of human behavior requires the agent to faithfully embody assigned roles and identities and proactively interact with other agents or humans (Franklin and Graesser, 1997; Elliott and Brzezinski, 1998; Laird and VanLent, 2001; Ortony et al., 2003; Wooldridge and Jennings, 1995; Jennings et al., 1998). Choi et al. (2023); Lin and Chen (2023); Fang et al. (2023); See et al. (2019) propose evaluations toward LLMs’ natural language understanding and generation. Rao et al. (2023b); Jiang et al. (2023b); Huang et al. (2023) evaluate LLMs’ ability to understand and maintain personality traits. Aher et al. (2023a) introduces the Turing Experiment to evaluate whether LLMs can simulate the behavior of a representative sample of participants in human subject research. Park et al. (2023); Gao et al. (2023a) propose a sandbox and an online social network to evaluate agents’ social interactions, respectively. However, to the best of our knowledge, no systematic and fine-grained benchmark exists to assess the agents’ believability. Hence, we aim to bridge this gap by proposing two novel metrics and constructing the first benchmark for simulation.
@@ 3 SimulateBench
To facilitate the measure of agents’ consistency and robustness, we introduce SimulateBench for characters’ profile collection and believability evaluation. We have collected the profiles of 56 characters and variants. There are 8400 questions in the consistency and robustness dataset. The statistic is shown in Table  1.
3.1 Profile Descriptive Framework and Datasets
Comprehensive profile information is necessary for agents to simulate human behavior accurately. Accordingly, we propose the profile descriptive framework and collect two character datasets based on this framework. For more details, please refer to the Appendix A.
@@ 3.1 Profile Descriptive Framework and Datasets
Researchers in different fields pay attention to distinct profile information, given their different objectives. We propose a descriptive framework that documents a character’s information from three aspects: Immutable Characteristic, Social Role, Relationship. Immutable characteristic refers to characteristics that cannot be easily changed, such as name, gender, and age. Social role (Wasserman and Faust, 1994; Eagly and Wood, 2012) is conceptualized as a set of connected behaviors, obligations, beliefs, and norms as conceptualized by people in a social situation.
Character Dataset
We selected characters from TV dramas of popular genres [^2] https://www.imdb.com/list/ls023983860/: The Simpsons (Animated), Friends (Comedy), Breaking Bad (Crime), and The Rings of Power(Science fiction). According to the profile descriptive framework, we extract the profile information from the fandom [^3] https://www.fandom.com/. Human annotators carefully verify the collected data to ensure that the profiles are error-free. We also collected real people’s profile data to enhance the diversity of the dataset. To ensure privacy protection, we blur sensitive information, including name, age, and location.
Character Variants Dataset
To facilitate the measurement of LLMs’ robustness, we perturb the profile of characters in the character dataset by replacing the content of demographic factors: Education, Surname, Race, and Age. To prevent irrationality caused by the perturbation, a thorough examination of the consequences resulting from any modifications made to the initial profile is conducted. For instance, if Homer’s educational background is altered, any mention of it will likewise be adjusted.
3.2 Consistency Dataset
The consistency dataset is composed of single-choice questions. The agent needs to answer the questions in the dataset. According to the profile descriptive framework, there are three kinds of questions related to Immutable Characteristics, Social Roles, and Relationships. We manually generate the gold answers for every question and double-check the correctness of the answer. For every kind of question, we further categorize the questions into two classes according to their gold answer: Known and Not-Known. The gold answer of Not-Known is “There’s not enough information to answer this question.”. The Not-Known is the case that there is not enough information about the character in the profile for the agent to infer the answer.
3.3 Robustness Dataset
The robustness dataset is collected based on the consistency dataset. We have perturbed the character’s profile for each character in the character dataset to construct the character variants dataset. According to this perturbation, we modify the corresponding questions in the consistency dataset. Then, we include the modified questions in our robustness dataset. For instance, if we modify the age of Homer from 20 to 30, our initial step will involve duplicating the questions pertaining to Homer in the consistency dataset. Subsequently, we shall alter these questions and their gold answer to align with the age adjustment. After the alteration of these questions, we get the questions for Homer at the age of 30. All these altered questions for character variants make up the robustness dataset.
@@ 3.3 Robustness Dataset
Table 1: The statistics of SimulateBench. The tokens are counted with the embedding of GPT-4.
4 Simulation and Measuring
The two key aspects of believability are consistency and robustness. To measure the consistency and robustness ability of agents, we have three steps: (1) employing the agents to simulate human behavior of assigned character by prompt engineering; (2) measuring the consistency by requiring the agents to answer the single-choice questions in the consistency dataset; and (3) measuring robustness by perturbing the profile of the assigned character and comparing how the agent’s consistency ability changed faced with the profile perturbation.
4.1 Human Behavior Simulation
To construct an LLM-based Agent, the LLM is usually used as the core brain to process reasoning, memory recall, and other cognitive activities. So, LLM is crucial for the agents’ human behavior simulation. Given that, We employ LLM as an agent (Liu et al., 2023c) without integrating other modules to focus on the LLMs’ function for simulation444we use LLMs and LLM-based agents interchangeably.. To prompt the LLM to simulate human behavior, some factors are crucial: the instruction to explain how to simulate human behavior (I), the profile of specific characters (P), and some description of the task that the LLM needs to finish (T). Formally, the LLMs ($P$) are supposed to generate a response $\text{Y}=y_{1},\ldots,y_{n}$:
[…]
Blew will introduce the setting of I, P, and T.

### v2

# How Far Are LLMs from Believable AI? A Benchmark for Evaluating the Believability of Human Behavior Simulation

@@ Abstract
In recent years, AI has demonstrated remarkable capabilities in simulating human behaviors, particularly those implemented with large language models (LLMs). However, due to the lack of systematic evaluation of LLMs’ simulated behaviors, the believability of LLMs among humans remains ambiguous, i.e., it is unclear which behaviors of LLMs are convincingly human-like and which need further improvements. In this work, we design SimulateBench to evaluate the believability of LLMs when simulating human behaviors. In specific, we evaluate the believability of LLMs based on two critical dimensions: 1) consistency: the extent to which LLMs can behave consistently with the given information of a human to simulate; and 2) robustness: the ability of LLMs’ simulated behaviors to remain robust when faced with perturbations. SimulateBench includes 65 character profiles and a total of 8,400 questions to examine LLMs’ simulated behaviors. Based on SimulateBench, we evaluate the performances of 10 widely used LLMs when simulating characters. The experimental results reveal that current LLMs struggle to align their behaviors with assigned characters and are vulnerable to perturbations in certain factors.  [^1] Code and SimulateBench are available at an anonymous GitHub repository.
How Far Are LLMs from Believable AI? A Benchmark for Evaluating the Believability of Human Behavior Simulation
Yang Xiao1  Yi Cheng1  Jinlan Fu2 Jiashuo Wang1Wenjie Li1thanks:    Corresponding AuthorsPengfei Liu3footnotemark: 1The Hong Kong Polytechnic University  2National University of Singapore 3Shanghai Jiao Tong University {yang-alan.xiao,alyssa.cheng}@connect.polyu.hkjinlanjonna@gmail.com csjwang@comp.polyu.edu.hkcswjli@comp.polyu.edu.hkpengfei@sjtu.edu.cn
@@ 1 Introduction
AI has shown promise to simulate human behavior and social interaction (Wooldridge and Jennings, 1995; Macal and North, 2005), which can empower applications ranging across prototyping social theories (Aher et al., 2023; Horton, 2023; Kovač et al., 2023), generating synthetic research data (Hämäläinen et al., 2023; Wang et al., 2023a) and building non-player characters (Laird and VanLent, 2001). These applications necessitate the simulated human behavior to possess a convincing level of believability, which allows the users to suspend their disbelief (Ortony et al., 2003). Such believability is crucial as it facilitates users in establishing trust in the AI and streamlines the fulfillment of the AI’s goals in these applications.
@@
Figure 1: An illustrative example of the “Consistency”, and “Robustness”. Consistency measures whether the LLMs’ generated human behavior accurately depicts the profile information; Robustness measures whether the generated human behavior will be influenced by the perturbation in the profile.
@@
Recently, the rapid development of LLM has accelerated the realization of human behavior simulation (Park et al., 2023), commonly achieved by providing the LLM with a relevant identity profile in the prompt. Despite the initial progress achieved by now, some challenging issues of LLMs pose a significant threat to the believability of LLMs: (1) Current LLMs cannot effectively process long input context (Liu et al., 2023). This could impede LLMs from capturing crucial information about characters in the profile, as the provided profile is usually long and complex, consisting of personal data such as identity information, social roles (Wasserman, 1994), and relationships (Hinde, 1976), to facilitate a comprehensive simulation. (2) LLMs lack robustness when confronted with perturbation to the input (Perez and Ribeiro, 2022). This could result in varying behaviors in the same scenarios. Especially when the environment where the LLMs are situated is dynamically changing, causing the profile information to be inevitably updated. LLMs’ believability would be compromised if they cannot effectively handle the constantly changing profile when simulating human behavior. Nevertheless, current studies (Park et al., 2023; Wang et al., 2023b; Shao et al., 2023; Cheng et al., 2023) fail to investigate the above two issues comprehensively. Firstly, these works only evaluate the LLMs’ capability to simulate very brief profile information, which is far from sufficient to simulate a person effectively. Moreover, there is little or no mechanism to thoroughly examine the potential impact of perturbations on the believability of LLMs and how these perturbations specifically affect their believability.
@@
To assess the negative impact brought by LLMs’ deficiency in processing long profile input and lack of robustness, we propose two dimensions to evaluate the believability of LLMs, as shown in Figure 1: (1) consistency: to what extent does the generated human behavior accurately depict the identity information, social roles, and relationships presented in the long profile input? (2) robustness: to what extent do the LLMs’ behaviors maintain robustness when faced with updates or perturbations in the profile? To measure consistency and robustness, we introduce SimulateBench, a benchmark for character data collection and evaluation of consistency and robustness. SimulateBench consists of four parts: the profile descriptive framework, the character profile dataset, the consistency dataset, and the robustness dataset. The profile descriptive framework is proposed to document the information about a person comprehensively. Based on the framework, we collect a character profile dataset, including the profiles of 65 characters. To measure the consistency, we assess whether the LLMs can correctly answer multi-choice questions about the character in the consistency dataset. To correctly answer these questions, the LLMs have to participate in logical reasoning based on the profile information. To measure the robustness, we perturb the profile of the assigned character and compare how the LLMs’ consistency ability changed based on the robustness dataset.
@@
Through the SimulateBench, we evaluate the level of believability of ten widely used LLMs. Our findings show that 1) LLMs perform poorly for consistency: they can not accurately depict the information in the long profile input, even if they are equipped with long context size; 2) LLMs exhibit a lack of robustness when faced with even nuanced profile perturbation; 3) LLMs exhibit similar poor robustness performance to different profile perturbations. In further studies, we examine four influential factors that will greatly influence the LLMs’ believability.
@@
In summary, we propose two novel dimensions of consistency and robustness to measure LLMs’ believability. To facilitate the assessment, we introduce the SimulateBench. We hope our work will inspire further research into the believability of human behavior simulation.
@@ 2 Related Work
2.1 Human behavior Simulation
Recently, LLMs have demonstrated intelligence comparable to humans in certain tasks (bench authors, 2023; Brown et al., 2020; Touvron et al., 2023). Many studies endeavor to harness the LLMs to simulate human behavior and social interactions in social science, economics, psychology, and human-computer interaction for prototyping theories and generating synthetic research data (Park et al., 2022, 2023; Argyle et al., 2023; Horton, 2023; Hämäläinen et al., 2023). Other studies prompt LMs(LLMs) with profiles to simulate human conversations in role-playing and personalized dialogue (Zhang et al., 2018; Zheng et al., 2019, 2020; Wang et al., 2023b). However, the profile of their work is concise. The limited amount of personal information provided is insufficient for the model to acquire sufficient knowledge to simulate a character accurately.
2.2 Evaluation of LLMs in Human Behavior Simulation
Simulation of human behavior requires the LLMs to faithfully embody assigned roles and identities and proactively interact with others (Wooldridge and Jennings, 1995; Franklin and Graesser, 1996; Ortony et al., 2003). See et al. (2019); Fang et al. (2023); Choi et al. (2023) propose evaluation frameworks toward LLMs’ capabilities of natural language understanding and generation. Rao et al. (2023); Jiang et al. (2023); Huang et al. (2023) evaluate LLMs’ abilities to understand and maintain personality traits. Aher et al. (2023) introduce the Turing Experiment to evaluate whether LLMs can simulate the behavior of a representative sample of participants in human subject research. Park et al. (2023) propose a sandbox and an online social network to evaluate agents’ social interactions. However, to the best of our knowledge, no systematic and fine-grained benchmark exists to assess the LLMs’ believability. Hence, we aim to bridge this gap by constructing SimulateBench.
@@ 3 SimulateBench
We introduce SimulateBench for character profile collection and believability evaluation. Specifically, our benchmark includes the profiles of 65 characters and 8400 questions to assess the LLMs’ consistency and robustness when simulating human behavior. The statistics are shown in Table  1.
3.1 Profile Descriptive Framework and Character Dataset
Comprehensive profile information is necessary for LLMs to simulate human behavior accurately. Accordingly, we propose the profile descriptive framework and collect a character dataset based on this framework. For more details, please refer to the Appendix A.
@@ 3.1 Profile Descriptive Framework and Character Dataset
We propose a descriptive framework that comprehensively documents a character’s information from three aspects: Immutable Characteristic, Social Role, Relationship. Immutable characteristic (Stein, 2001) refers to characteristics that cannot be easily changed, such as name, gender, and age. Social role (Wasserman, 1994; Eagly and Wood, 2012) is conceptualized as a set of connected behaviors, obligations, beliefs, and norms as conceptualized by people in a social situation. Relationship (Sztompka, 2002) is the basic element of study in the field of social sciences and refers to any interpersonal connection between two or more individuals. Furthermore, these three kinds of profile information are thoroughly elaborated by fine-grained aspects based on established theories. For example, we will comprehensively document the following aspects of the relationship: familiarity, judgment, affection, behavioral patterns, relationship status, and communication history. For more details, please refer to the Appendix A.1.
@@
Table 1: The statistics of SimulateBench. The tokens are counted with the tokenizer of GPT-4.
Character Dataset
We selected characters from TV dramas of popular genres [^2] https://www.imdb.com/list/ls023983860/: The Simpsons (Animated), Friends (Comedy), Breaking Bad (Crime), and The Rings of Power (Science fiction). According to the profile descriptive framework, we extract the profile information from the fandom [^3] https://www.fandom.com/. As shown in Table  1, the profile in our work contains 3,277 tokens on average, which is comprehensive in comparison to prior studies. As an illustration, the profile mentioned in the well-known study by Park et al. (2023) only contains 203 tokens.
We recruit four human annotators to carefully verify the collected data and ensure the profiles are error-free. Each annotator completed an extensive ten-hour training program that covered the functioning of the profile descriptive framework and the methods for detecting data errors. If they encounter characters that they find unappealing or unfamiliar, we will halt their annotation process.
3.2 Measuring Consistency
Consistency Dataset
The consistency dataset is composed of multi-choice questions. Each character has an average of 150 questions. To accurately answer these questions, the LLMs needs to analyze and employ logical reasoning to the profile information. We apply GPT-4 to autonomously generate these questions based on the various profile information types. For question generation details and question examples, please refer to the Appendix A.4. According to the profile descriptive framework, there are three kinds of questions related to Immutable Characteristics, Social Roles, and Relationships. We manually generate the gold answers for every question and double-check the correctness of the answer.
For every question, we add a choice of “There’s not enough information to answer this question”. This choice is intended for the case that there is no sufficient information about the character in the profile for the LLM to deduce the answer, and we set this choice as the gold answer in such a case. The reason for this setting is that if the LLM is given unrestricted freedom to respond to the content that is not mentioned in the profile, there is a high probability of compromising the character’s information and undermining the LLM’s believability. We further categorize the questions into two classes according to their gold answer: Known and Unknown. The gold answer of Unkown is “There’s not enough information to answer this question”. For question examples, please refer to the Appendix A.4.
Measuring Metric: CA
To measure the consistency, we will employ the LLMs to answer the questions in the consistency dataset, and we will calculate the accuracy of these answers as the consistency ability, referred to as CA.
3.3 Measuring Robustness
Robustness Dataset
The robustness dataset is constructed by conducting perturbations on the characters’ profiles (denoted by the characters’ variant) and modifying the questions in the consistency dataset accordingly. We perturb the profile of characters by replacing the content of demographic factors: Education, Surname, Race, and Age. To prevent irrationality caused by the perturbation, a thorough examination of the consequences resulting from any modifications made to the initial profile is conducted. Please refer to the Appendix A.3 for more details. According to this perturbation, we modify the corresponding questions in the consistency dataset. Then, we include the modified questions in our robustness dataset. For instance, if we modify the age of a character from 20 to 30, our initial step will involve duplicating the questions pertaining to the character in the consistency dataset. Subsequently, we shall alter these questions and their gold answers to align with the age adjustment. After the alteration of these questions, we get the questions for the character at the age of 30.
Measuring Metrics: RA and RCoV
The robustness aims to determine the variation in the consistency performance of the LLMs when slight modifications are made to profiles. To achieve this goal, we employ the variation of CA and coefficient of variation [^4] https://en.wikipedia.org/wiki/Coefficient_of_variation of CA scores as the robustness performance of LLMs, referred to as RA and RCoV respectively. For example, when employing GPT-3.5 to simulate a character, only modifying the age property in the profile to values of 10, 15, 20, 25, and 30 yields five variants. After all five variants finish the robustness dataset, five CA scores will exist: $s_{1},\ldots,s_{5}$. The five scores’ deviation and mean are $\sigma$, $\mu$. The RA of GPT-3.5 will be $RA=\sigma$. The RCoV of GPT-3.5 will be $RCoV=RA/\mu$.
Dividing RA by $\mu$ allows for the comparison of different models. RCoV can be understood as the quantification of the impact that robustness (RA) can have on the actual performance ($\mu$). As an illustration, LLM A demonstrates an RA of 0.04, a $\mu$ of 0.3, and hence RCoV to be 0.13. LLM B exhibits an RA of 0.08, a $\mu$ of 0.9, and hence RCoV to be 0.089. While LLM B has a higher RA score (0.08 compared to 0.04), the actual impact of RA on performance is smaller (0.089 compared to 0.13).
Immutable Characteristic Social Role Relationship
Model CA Known Unknown Known Unknow Known UnKnown
GPT-4 0.77 1.00 0.47 1.00 0.59 0.97 0.06
GPT-3.5-Turbo-16K 0.70 0.82 0.58 0.56 0.88 0.91 0.31
XVERSE-13B-Chat 0.62 0.68 0.53 0.68 0.76 0.59 0.44
Qwen-14B-Chat 0.60 0.59 0.21 0.82 0.12 0.94 0.38
Vicuna-13B-16K 0.59 0.64 0.32 0.76 0.18 0.76 0.56
ChatGLM2-6B-32K 0.55 0.68 0.21 0.71 0.24 0.79 0.25
Qwen-7B-Chat 0.53 0.64 0.11 0.74 0.12 0.91 0.06
ChatGLM2-6B 0.49 0.50 0.16 0.65 0.12 0.88 0.06
LongChat-7B-32K 0.48 0.59 0.05 0.76 0.00 0.79 0.06
Vicuna-7B-16K 0.46 0.36 0.05 0.85 0.06 0.74 0.06
Table 2: CA scores across ten models to simulate a character. The last six columns correspond to the accuracy of the model for different types of questions. A larger CA indicates better consistency performance.
4 Baseline Methods for Human Behavior Simulation
To prompt the LLM to simulate human behavior, three components are crucial: the instruction to explain how to simulate human behavior (I), the profile of specific characters (II), and some description of the task that the LLM needs to finish (III). Below will introduce how we implement these three components in our baselines.